LASER INDUCED FLUORESCENCE AND LIFETIME MEASUREMENTS OF ORGANIC RADICALS

\underline{S.A. WILLIAMS}, E. ZINGHER, and J.C. WEISSHAAR, Department of Chemistry, University of Wisconsin - Madison, Madison, WI 53706.

Different isomers of vinoxy radical were produced in a pulsed supersonic expansion by excimer laser photolysis at 193 nm. The B̃ to X̃ electronic transitions were probed by laser induced fluorescence. The spectra are complicated. We find as many as 30 bands over a 2700 cm^{−1} interval. We have measured fluorescence lifetimes of many of these bands. The lifetimes decrease slowly toward higher energies to as short as 25 ns and then the bands abruptly end. Comparisons with ab initio work will be discussed.